Using the method of undetermined coefficients, find the general solution of the differential equation,
y'''' - 2y''' + 2y'' = 3e^(-x) + 2e^(-x)x + e^(-x)sinx.
1
Expert's answer
2020-04-29T18:04:44-0400

"y^{IV}-2y^{III}+2y^{II}=(3+2x)e^{-x}+e^{-x} \sin x"

Solution find in form

"y=y_0+y_1+y_2"

1) "y_0" find from equation

"y^{IV}-2y^{III}+2y^{II}=0\\
\lambda ^4-2\lambda^3+2\lambda^2=0\\
\lambda_1=\lambda_2=0, \lambda_3=1-i, \lambda_4=1+i\\
y_0=C_1+C_2x+C_3e^x\cos x+C_4e^x \sin x"


2) "y_1" find in form

"y_1=(ax+b)e^{-x}"

for equation

"y^{IV}-2y^{III}+2y^{II}=(3+2x)e^{-x}\\
y_1'=(a-ax-b)e^{-x}\\
y_1''=(ax-2a+b)e^{-x}\\
y_1'''=(3a-ax-b)e^{-x}\\
y_1''''=(ax-4a+b)e^{-x}\\"

input in equation

"(ax-4a+b)e^{-x}-2(3a-ax-b)e^{-x}+\\
+2(ax-2a+b)e^{-x}=(3+2x)e^{-x}\\
x:a+2a+2a=2\\
x^0:b-4a-6a+2b-4a+2b=3\\
a=\frac{2}{5}\\
b=\frac{43}{25}\\
y_1=(\frac{2}{5}x+\frac{43}{25})e^{-x}"


3) "y_2" find in form

"y_2=ke^{-x}\cos x+me^{-x}\sin x"

for equation

"y^{IV}-2y'''+2y''=e^{-x}\sin x\\
y'_2=(m-k)e^{-x}\cos x-(k+m)e^{-x}\sin x\\
y''_2=-2me^{-x}\cos x+2ke^{-x}\sin x\\
y'''_2=(2m+2k)e^{-x}\cos x+(2m-2k)e^{-x}\sin x\\
y''''_2=-4ke^{-x}\cos x-4me^{-x}\sin x\\
-4ke^{-x}\cos x-4me^{-x}\sin x-\\
-2(2m+2k)e^{-x}\cos x-2(2m-2k)e^{-x}\sin x-\\
-4me^{-x}\cos x+4ke^{-x}\sin x=e^{-x}\sin x\\
\cos x: -4k-4m-4k-4m=0\\
\sin x:-4m-4m+4k+4k=1\\
m=-\frac{1}{16}\\
k=\frac{1}{16}\\
y_2=\frac{1}{16}e^{-x}\cos x-\frac{1}{16}e^{-x}\sin x"

than

"y=C_1+C_2x+C_3e^x\cos x+C_4e^x\sin x+\\
+(\frac{2}{5}x+\frac{43}{25})e^{-x}+\frac{1}{16}e^{-x}\cos x-\frac{1}{16}e^{-x}\sin x"